5 . Problems and Applications Q1 A large share of the world supply of diamonds comes from Russia and South Africa, Suppose that the marginal cost of mining diamonds is constant at $3,000 per diamond, and the demand for diamonds is described by the following schedule Price Quantity (Dollars) (Diamonds) 8,000 7,000 6,000 5,000 4,000 3,000 2,000 1,000 10,000 3,000 4,000 5,000 6,000 7,000 8,000 9,000
Answer #1
Price Quantity TR MR
(Dollars) (Diamonds)
8,000 3,000 24000000
7,000 4,000 28000000   4000
6,000 5,000 30000000 2000
5,000 6,000 30000000 0
4,000 7,000 28000000 -2000
3,000 8,000 24000000 -4000
2,000 9,000 18000000 -6000
1,000 10,000 10000000 -8000

If there were many suppliers of diamonds, the price would be P=MC = 3000 per diamond and the quantity sold would be 8000 diamonds.

If there were only one supplier of diamonds, the price would be MC=MR =7000 per diamond and the quantity sold would be 4000 diamonds.

Suppose Russia and South Africa form a cartel.

In this case, the price would be 7000 per diamond and the total quantity sold would be 4000 diamonds. If the countries split the market evenly, South Africa would produce 4000/2 =2000 diamonds and earn a profit of 2000*(7000-3000) = 8000000

If South Africa increased its production by 1,000 diamonds while Russia stuck to the cartel agreement, South Africa's profit would [increase] to 3000*(6000-3000) = 9000000

Why are cartel agreements often not successful?.

B) One party has an incentive to cheat to make more profit.